### v1

# Multimodal Financial Foundation Models (MFFMs): Progress, Prospects, and Challenges

Abstract.
Financial Large Language Models (FinLLMs), such as open FinGPT and proprietary BloombergGPT, have demonstrated great potential in select areas of financial services. Beyond this earlier language-centric approach, Multimodal Financial Foundation Models (MFFMs) can digest interleaved multimodal financial data, including fundamental data, market data, data analytics, macroeconomic, and alternative data (e.g., natural language, audio, images, and video). In this position paper, presented at the MFFM Workshop joined with ACM International Conference on AI in Finance (ICAIF) 2024, we describe the progress, prospects, and challenges of MFFMs. This paper also highlights ongoing research on FinAgents in the SecureFinAI Lab [^1] https://openfin.engineering.columbia.edu/ at Columbia University. We believe that MFFMs will enable a deeper understanding of the underlying complexity associated with numerous financial tasks and data, streamlining the operation of financial services and investment processes.
Github Repo: https://github.com/Open-Finance-Lab/Awesome-MFFMs/
copyright: acmlicensedjournalyear: 2024copyright: acmlicensedconference: International Workshop on Multimodal Financial Foundation Models (MFFMs) at 5th ACM International Conference on AI in Finance; Nov. 15; NY, USAbooktitle: International Workshop on Multimodal Financial Foundation Models (MFFMs) at 5th ACM International Conference on AI in Finance (MFFM at ICAIF ’24), November 14–17, 2024, Brooklyn, NY, USA
Figure 1. An overview. Multimodal FinData (left block) is ubiquitous in FinApps (middle block), such as search, tutor, robo-advisor, compliance, auditing, trading, etc. However, several major challenges (right block) call for immediate actions, in order to achieve FinAI readiness and governance.
1. Introduction
The general public could not afford a private lunch with billionaire Warren Buffett. What about hiring Buffett as my on-call financial advisor, possibly at a cost of \$100? Before finalizing an investment of \$50 million, how about holding an elite brainstorming session with the world’s 10 greatest investors?…at a cost of \$5000?… Not many years later, as he earned his first \$1 million from investment, Yanglet was to remember that night when Dr. Li Deng helped him revise the proposal for the Multimodal Financial Foundation Models (MFFMs) workshop.
Large language models (LLMs) have demonstrated remarkable proficiency in understanding and generating human-like texts. FinLLMs, such as open FinGPT (Liu et al., 2023, 2024a; Yang et al., 2023; Zhang et al., 2023a; Tian et al., 2024) and proprietary BloombergGPT (Wu et al., 2023), have shown great potential in select areas of financial services (Kong et al., 2024a, b) . Beyond this earlier language-centric approach, Multimodal Financial Foundation Models (MFFMs) can process interleaved multimodal financial data, including fundamental data, market data, data analytics, macroeconomics, and alternative data (e.g. natural language, audio, image and video). Multimodal financial data has unique characteristics, such as dynamic, both structured and unstructured forms, and comes in varying formats (e.g., charts, graphs, Web APIs, Excel spreadsheets, SEC filings, XBRL filings (Han et al., 2024), and SQL data (Zhao et al., 2024)).
On the journey toward widespread adoption of MFFMs, several challenges remain, including increasing concerns related to reproducibility, transparency, privacy, and ethics. First, many existing LLMs function as black boxes, posing challenges in comprehending their operations and ensuring fairness. Another two major challenges are “model cannibalism” and “openwashing.” Many models are largely trained and released without transparency in mind, e.g., Claude 3.5 Sonnet. Many supposedly ”novel” models may exploit labels from existing LLMs (e.g., GPT-4o) and perform supervised learning, referred to as “model cannibalism.” As a result, MFFMs are opaque in decision-making. They give rise to various challenges, such as inadequate transparency concerning the training data, deficiencies in combating models’ inherent biases, safety and security issues, and adversarial attacks (e.g., backdoor attacks).
Recently, there have been many “openwashing” behaviors, where the LLM weights are marketed as ”open” but under restricted licenses, while OSI-approved licenses (e.g., Apache License 2.0 and MIT License) are preferable. A research alliance among Columbia University, Oxford University, and the Linux Foundation proposed the Model Openness Framework (White et al., 2024) that classifies model openness by ranking models according to $1616$ components. This framework offers guidance to researchers and model producers for promoting transparency and reproducibility. MFFMs that comply with this framework would promote reproducibility and adoption across the finance industry, such as Open FinLLMs (Xie et al., 2024b). For financial institutions, it provides clear guidelines for new models to become commercially suitable without restrictions.
In this position paper, presented at the MFFM Workshop [^2] MFFM Website: https://sites.google.com/view/iwmffm2024/home?authuser=1 jointly held with ACM International Conference on AI in Finance (ICAIF) 2024, we describe the progress, prospects, and challenges of MFFMs. This paper also highlights ongoing research on FinAgents in the SecureFinAI Lab at Columbia University. We first list multimodal financial data and data-centric approach in the financial domain (left block of Fig. 1). Then, we describe multimodal financial applications (middle block of Fig. 1). We envision that the AI agent is a promising solution to build multiple financial applications. Using the Model Openness Framework, licenses (e.g. OpenMDW [^3] https://openmdw.ai/) and agent protocols (e.g. Model Context Protocol, Agent2Agent Protocol) can build financial Agentic AI ecosystems. However, several major challenges (right block of Fig. 1) call for immediate actions in order to achieve financial AI readiness. The associated challenges are proprietary data constraints, training and inference costs, regulatory complexities, reasoning capacity, and the need for robust benchmarks and a guardrail framework to address misinformation and data biases. We believe that MFFMs will enable promising financial tasks and data analysis, streamlining the operation of financial services.
Survey Date FinLLMs Benchmark Applications Challenges Multimodal Readiness/Governance
Li et al. (2023) Nov. 2023
Ding et al. (2024) Jul. 2024
Lee et al. (2024) Apr. 2024
Kong et al. (2024a, b) Late 2024
This Survey Jan. 2025
Table 1. Overview of related surveys. The square indicates that the topic was covered but not comprehensive.
Related Work: There are several related surveys on FinLLMs, as outlined in Table 1. Li et al. (2023) first reviewed the approach employing LLMs in finance. Ding et al. (2024) summarized the performance of LLM-based agents in financial trading tasks. Lee et al. (2024) reviewed FinLLMs from a benchmark perspective. Kong et al. (2024a) and Kong et al. (2024b) further summarize recent advancements in FinLLMs and discuss their various application scenarios. Despite these efforts, these surveys lack a comprehensive review of multimodal financial data, applications, and models. Furthermore, there is still a need for an in-depth analysis concerning the opportunities and challenges of applying LLMs in finance, including their current readiness level.
Contributions: We summarize the contributions as follows:
  • To the best of our knowledge, this is the first comprehensive survey of multimodal financial foundation models (MFFMs). We summarize three aspects, multimodal financial data, applications, and model exploration.
  • For multimodal financial data, we emphasize a data-centric approach. For applications, we point out the trend of agentic AI ecosystem, which is enabled by open models and the model context protocol.
  • We compare and contrast MFFMs with LLMs, FinLLMs, and MM-LLMs. Our aim is to offer readers a holistic view of MFFM development and help readers understand the current progress and future prospects.
  • We describe the opportunities and point out the challenges when applying MFFMs in the financial domain, including proprietary data and digital regulatory reporting.
  • We discuss the ethical challenges of MFFMs’s readiness, including the hallucination and misinformation that may arise from real-world use scenarios, as well as ethical issues.
Well, using an MFFM base model, one could fine-tune a “Warren Buffett” model using the QLoRA method as in (Liu et al., 2023, 2024a) by feeding multimodal data, e.g., Buffett‘s conference transcripts, audio, video, interviews, and the fine-tuning cost would be less than \$100. On the other hand, by specifying the preferred articles/websites or creating a customized database as in (Tian et al., 2024), an investment institution would consult the world’s ten greatest investors (namely their digital avatars) in an elite brainstorming session at the cost of \$5000.
Key Terms Explanations
Transformer A transformer is a neural network architecture that utilizes the multi-head attention mechanism.
Large Language Model (LLM) LLM is a type of machine learning model for human-like text understanding and generation.
Pre-training Pre-training refers to the initial training phase where a model learns general features from a large dataset.
Fine-tuning Since a pre-trained LLM has a large number of parameters, trained on a huge dataset over millions of GPU hours, it is natural to employ a fine-tuning method to scale such a GPT model to hundreds of use scenarios.
Generative Pre-trained Transformer (GPT) GPT is a family of LLMs based on a transformer architecture.
Prompt engineering The process of structuring an instruction in order to produce the best possible output from an LLM model.
Zero-Shot Prompting An LLM is given a task without examples or training on that task, relying on LLM’s pre-existing knowledge to generate a response.
Few-Shot Prompting The prompt of providing a generative model with a few examples of a task to guide its output.
Chain-of-Thoughts (CoT) A prompt engineering strategy to guide language models to handle complex reasoning tasks. For example, write the reasoning guidance in the prompt.
In-Context Learning (ICL) ICL is a new learning paradigm where a language model observes a few examples and directly outputs the test input’s prediction.
Foundation Model A foundation model is a machine learning or deep learning model that is trained on vast datasets so it can be applied across a wide range of downstream tasks.
FinLLM A foundation model for financial applications.
Multimodal Multimodal means “having several modalities”, and a ”modality” refers to a type of input or output, such as video, image, audio, text, proprioception, etc.
Retrieval-Augmented Generation (RAG) RAG is a process of optimizing the output of an LLM. It references an authoritative knowledge base outside of its training data sources before generating a response.
Low-Rank Adaptation (LoRA) LoRA is a popular and efficient training technique that significantly reduces the number of trainable parameters.
QLoRA QLoRA is the extended version of LoRA, which works by quantizing the precision of the weight parameters in the pre-trained LLM to 4-bit precision.
Agent A decision maker. The LLM-powered agent is a powerful framework for solving complex tasks by using an LLM as its central computational engine.
Model Context Protocol (MCP) A standard for connecting AI assistants to the systems where data lives, including content repositories, business tools, and development environments.
Agent2Agent (A2A) Protocol A standard to facilitate communication between independent AI agents.
Openwashing A term is used to describe presenting some models as open source when they are not using permissive licenses.
Table 2. Terminology for LLMs and agents.
Key Terms Explanations
Earnings Conference Calls (ECCs) A call between a public company and key stakeholders to discuss the company’s financial results.
Monetary Policy Calls (MPCs) Countries’ central banks hold MPC to decide what monetary policy action to take.
Environmental, Social, Governance (ESG) This is shorthand for an investing principle that prioritizes environmental issues, social issues, and corporate governance.
Financial Decision Making It encompasses evaluating options, making choices, and taking actions (trading) related to financial matters.
eXtensible Business Reporting Language (XBRL) XBRL is the global standard that powers digital reporting.
Common Domain Model (CDM) CDM is a standardized, machine-readable, and machine-executable data and process model for how financial products are traded and managed across the transaction lifecycle.
Robo-Advisor A type of automated financial advisor that provides algorithm-driven management services without human intervention.
Digital Regulatory Reporting (DRR) DRR is a cross-industry initiative to transform the reporting infrastructure.
Greenwashing Promotes false solutions to the climate crisis that distract from and delay concrete and credible action.
Table 3. Terminology for finance.
2. Terminology
Multimodal Financial Foundation Models (MFFMs) is an intersection field of foundation models and finance. To facilitate readers from various backgrounds, we first provide two lists of terminologies in Table 2 and Table 3, respectively.
Types Text Audio Image Video Numbers Tabular Chart Time-Series
Earnings Conference Calls (ECC)
Monetary Policy Calls (MPC)
Climate Data
Financial News
Market Data
Financial Reports
Financial curriculum and certificates
Table 4. Overview of multimodal financial data.
3. Data-Centric Approach for Multimodal Financial Data
First, we summarize the common multimodal financial data in Section 3.1. Second, we present detailed descriptions of typical multimodal financial data types in Sections 3.2 to 3.7.
3.1. Spectrum of Multimodal Financial Data
Multimodal data is common in business, finance, accounting, and auditing, as illustrated in Fig. 1 (left block):
  • Textual data: Text is the most prevalent data type, including financial news, financial reports, earnings conference call transcripts, and social media posts. These textual data provide timely market information and reflect market participants’ sentiments.
  • Numerical data: Numerical data, such as stock prices, financial indicators, and economic statistics, offer market insights. Investors and analysts frequently rely on numerical data for market forecasting.
  • Chart data: Charts are frequently included in financial reports, news articles, and related materials. It visually represents market trends and patterns, facilitating easier interpretation of market behavior and dynamics.
  • Tabular data: Structured financial data presented in tables, including balance sheets, income statements, stock prices, and trading volumes.
  • Time-series data: It is a sequence of data points indexed in time order. In the financial sector, time series data is commonly used to represent how a financial indicator changes over time.
  • Visual data: Visual data includes images and videos. They are from financial media and official announcements. Visual data provide detailed insights beyond textual and numerical data, illustrating complex market events and trends.
  • Audio data: Financial podcasts and recordings of earnings conference calls contain critical auditory information. Audio modalities can influence market perception and offer additional dimensions for sentiment analysis and market prediction.
Multimodal financial data can refer to a combination of the above uni-modal data. For instance, Earning Conference Calls (ECCs) consist of two modalities: the audio of a presentation and its textual transcripts. We list the common types in Table 4 and describe them in the following subsections.
3.2. Earning Conference Calls (ECCs)
Figure 2. An ECC example of Tesla 2024 Q3 on Oct. 23, 2024. The CEO, Elon Musk, presented a speech to explain the company’s revenue for the past quarter and major related events and provided an overview plan. The close prices of the following two days were \$260.48currency-dollar260.48\\$260.48\$ 260.48 and \$269.19currency-dollar269.19\\$269.19\$ 269.19.
The earnings conference call (ECC) is a teleconference or webcast held quarterly by a public company. Stakeholders (including analysts, investors, and the media) participate to obtain the company’s latest financial status. First, the company’s CEO/CFO highlights the quarterly financial status, strategic initiatives, and forward-looking plans. Then, analysts and investors ask questions in the Q&A sessions. The release of ECCs is correlated with market reactions, making them an important resource for analyzing market changes (Foster et al., 1984).
An example of Tesla 2024 Q3 ECC, as shown in Fig.  2, has 72 minutes. The call includes 29 minutes presentation by Tesla CEO Elon Musk, followed by 43 minutes Q&A session. First, CEO Elon Musk summarized Tesla’s Q3 revenue and car production status and underscored Tesla’s ongoing strategy to hasten the global transition to sustainable energy. In the end, Musk reiterated Tesla’s preparations for introducing more affordable models. In the Q&A session, Tesla’s executive team responded to questions about product research and development, upcoming product plans, Tesla’s Full Self-Driving offerings, etc. Owing to good revenue performance and car production, Tesla’s stock price sustained an upward trend in the following two days. The entire ECC is saved as a ”.mp3/.wav” audio file, and the corresponding transcript is also recorded. Both audio and text data can be accessed or analyzed by the public.
The creation of an ECC dataset is critical for developing analytical tools, particularly for stock movement prediction and risk modeling. MDRM (Qin and Yang, 2019) is a representative ECC dataset that includes 576 earnings conference calls from 280 companies in the S&P 500 for the year 2017. The entire dataset is 5.7 GB in storage. The author segmented the transcripts into individual sentences and aligned them with corresponding audio clips, resulting in a total of 88,829 paired sentences and audio clips. The audio data is available from earningcast [^4] https://earningscast.com/ and transcript file can be downloaded from Seeking Alpha [^5] https://seekingalpha.com/. A series of research works extract critical information from textual transcripts and integrate it with audio features such as tone and sentiment in a speech to assess risks (e.g., volatility) (Qin and Yang, 2019; Yang et al., 2020; Cao et al., 2024a).
The current approach to financial analysis based on ECC data faces several key challenges related to dataset curation. First, the existing ECC dataset is limited in size and lacks sufficient coverage of companies across diverse industries, as ECC characteristics vary considerably between companies. Second, aligning audio with the text remains imperfect. Splitting the ECC data into segments frequently fails to align precisely with sentence boundaries, potentially leading to semantic incoherence in the segmented text and audio. Therefore, it is essential to establish a dataset curation pipeline that focuses on acquiring, organizing, segmenting, and labeling ECC data. Such a data infrastructure will enable the creation of more effective financial applications.
3.3. Monetary Policy Conferences (MPCs)
Monetary Policy Conferences (MPCs) are regularly held by a country’s central bank, like the U.S. Federal Reserve. An MPC deliberates on a nation’s economic conditions, articulates monetary policy, and assesses potential economic risks. The conference includes a press presentation by the governor, followed by a Q&A session with journalists. Given their provision of critical insights into a central bank’s decisions, MPCs are instrumental in influencing economic conditions, such as commodity trading, inflation, and currency exchange rates.
Figure 3. An example of Monetary Policy Conference held by the U.S. Federal Reserve. The Governor first presents a press speech, followed by a spontaneous Q&A session. After the conference, the stock index declined, while Treasury yields increased. In the period following the conference, the stock index exhibited a gradual recovery.
An example is the Federal Open Market Committee (FOMC) conference on May 1st, 2024, as shown in Fig.  3. This conference has 48 minutes with 8 minutes presentation and 40 minutes Q&A session. In the presentation, Federal Reserve Chairman Jerome Powell emphasized that inflation remains a major concern and suggested the possibility of further policy tightening. Then, the FOMC unanimously decided to keep the benchmark short-term borrowing rate at 5.25%–5.5%. This rate is the highest level in 23 years. After the release of the conference recording, consumer surveys indicated heightened economic anxiety [^6] Reported by CNBC: https://www.cnbc.com/2024/05/22/fed-minutes-may-2024-.html. Stock markets experienced declines while Treasury yields increased. In the subsequent period, however, the stock index gradually recovered.
Analyzing MPC data can help illuminate policy decisions and assist in forecasting economic trends. The first comprehensive MPC dataset MONOPOLY (Mathur et al., 2022) has 180 GB in size. It includes 340 MPC instances from six countries’ central banks: the United States, the United Kingdom, the European Union, Canada, New Zealand, and South Africa. In total, the dataset comprises 15,729 minutes of recorded content, where each MPC session has on average 53 minutes. Typically, every MPC includes approximately 10 minutes of presentation, followed by a Q&A session with more than 40 minutes. Each MPC consists of three parts: Audio, Text, and Video. The dataset was constructed by employing the BeautifulSoup Python package [^7] https://www.crummy.com/software/BeautifulSoup/ to scrape MPC dates, ‘.mp3’ audio, ‘.MP4’ videos, and PDF transcripts. Text data are subsequently extracted from PDF transcripts by using Urllib [^8] https://pypi.org/project/urllib3/. Recent studies use this dataset to jointly model audio, text, and video features to predict various economic indicators (Mathur et al., 2022; Ouyang et al., 2024).
The following challenges are faced when using the MPC dataset to build an analysis tool for economic conditions. First, storing MPC data needs to maintain audio, text, and video modalities together. It introduces the challenges related to efficient data management. Second, processing MPC data requires precise alignment across audio, text, and video modalities, which remains technically challenging. Thus, establishing a data curation pipeline for MPC is essential for advancing related methodological development.
Financial Reports Frequency Publisher Focus Required by SEC
Form 10-Q Quarterly Company Interim financial statements and recent operational updates Yes
Form 10-K Annually Company Comprehensive, audited financial results and business overview Yes
DEF 14A Annually Company Governance, executive compensation, and voting matters Yes
Form 8-K Event-driven Company Disclosure of significant, material events affecting operations Yes
Earnings Release Quarterly Company Preliminary quarterly financial results and management commentary No
Annual Report Annually Company Simplified summary highlighting performance and management vision No
Zacks Investment Reports Frequently Third-party (Zacks) Stock ratings, earnings forecasts, and investment recommendations No
Sell-side Broker Reports Frequently Third-party (Analysts) In-depth analysis, valuation, forecasts, and buy/sell recommendations No
Table 5. Financial reports with different frequency, source, primary focus and SEC requirements.
3.4. Financial Reports
The financial report is a formal document that presents a company’s financial activities, performance, management discussion, and audited financial statements. The frequently-used financial reports include filings (e.g., 10-K, 10-Q, DEF-14A, 8-K) required by the U.S. Securities and Exchange Commission (SEC), company-issued documents for stakeholders (e.g., earnings releases and annual reports), and third-party analysis reports such as Zacks reports and sell-side broker reports. These reports differ in their publication frequency, publisher, and areas of emphasis. A summary of these financial reports is provided in Table 5. Market participants can access various financial reports from different companies based on their specific needs. These reports enable investors to evaluate a company’s status and identify broader market trends. Additionally, these financial reports are monitored by government and regulatory agencies to ensure fairness in trading and other financial activities.
3.5. Financial News
Financial news refers to news that pertains to money and investments, including news on markets. It is disseminated through various channels, including traditional financial reporting (e.g., The Wall Street Journal), financial news services platforms (e.g., Bloomberg terminal), social media (e.g., Twitter and LinkedIn), online discussion forums (e.g., Reddit), and interactive media formats such as live broadcasts. Financial news can take different formats, including text, video, audio, numerical, charts, and tabular data. It has become increasingly important in forecasting financial outcomes, such as stock volatility, investor sentiment, market risks, and macroeconomic stability (Schumaker et al., 2012; Piškorec et al., 2014).
The GameStop (GME) short squeeze event in January 2021 exemplifies how financial news can impact the financial markets. In the beginning, hedge funds published short-selling reports on GME, forecasting a decline in its stock price based on weak financial fundamentals. They also spread their view online, prompting institutional investors to establish short positions on GME. However, individual investors on social platforms such as Reddit’s r/WallStreetBets rapidly disseminated relevant financial news and coordinated buying activities. Individual investors’ behavior drives GME stock prices upward. Additionally, Elon Musk’s retweet of related news further encouraged retail investor participation, amplifying the stock’s upward. Institutional investors subsequently faced pressure to purchase shares to cover their short positions, thereby intensifying the stock’s upward momentum. Within a short period, the GME stock price surged nearly 190 times, resulting in significant losses for hedge funds and substantial gains for retail investors. This incident shows the potent influence of financial news on market trends. Lin et al. (2024b) employed LLMs to analyze this phenomenon. LLMs were used to clean extensive collections of online financial news, resulting in the creation of a high-quality dataset. The dataset was then used to analyze user behavior and the underlying mechanisms driving information dissemination. This research emphasizes the significant role of financial news in the impact of financial markets.
Effective collection of financial news data is crucial for analyzing market dynamics. Financial news data can be collected from various online platforms and sources. First, specialized financial platforms, such as Bloomberg, Dow Jones, Yahoo Finance, and CNBC, deliver timely and professional financial news. Second, professional news organizations, such as Reuters, offer financial news coverage. Third, social media, including X (formerly Twitter) and Reddit, also provide financial news. Users can access financial news data through platform-specific APIs or manually gather financial news data directly from these platforms. However, users must be mindful of copyright restrictions associated with each platform.
Although amounts of financial news data are available, there remain several challenges to making these raw data into usable datasets: 1) Trustworthiness. Financial news from various sources may include subjective content or misinformation. Evaluating the reliability of financial news is a significant challenge. 2) Volume issue. A large amount of financial news is disseminated daily, making it difficult to effectively process and manage. 3) Modality alignment. Financial news includes various types of information, such as charts, tables, and images. A key challenge is accurately aligning textual content with its corresponding other elements.
3.6. Market Data and Alternative Data
3.6.1. Market Data
Market Data refers to price information and other related data for financial instruments provided by trading venues. It represents financial information through different modalities. For example, market data uses time-series data to record a company’s stock price, numerical data to represent financial indicators, and charts or tabular data to display a company’s operational performance. These multimodal market data provide investors with diverse perspectives on current market changes and historical movements, supporting informed decision-making (Lee and Yoo, 2020).
Financial markets have undergone a rapid change due to the increasing amount of data. Extracting actionable insights from these vast and heterogeneous market data to support decision-making in complex market environments has therefore become a challenge (Hambly et al., 2023). Reinforcement learning (RL) provides a promising approach to solve this challenge. RL could train trading agents to interact with dynamic market environments and to optimize their financial decisions autonomously (Hambly et al., 2023; Sun et al., 2023). The Financial reinforcement learning (FinRL) project (Liu et al., 2018, 2020, 2021, 2022) offers a user-friendly virtual market environment that includes a wide range of multimodal market data. FinRL integrates commonly used Deep Reinforcement Learning (DRL) algorithms, enabling users to develop their own DRL trading strategies. Recently, FinRL 2025 contest [^9] https://finrl-contest.readthedocs.io/en/latest/ proposed the FinRL-DeepSeek project, combining reinforcement learning with LLMs to develop an automated stock trading agent trained on stock price and financial news data. This hybrid approach enhances the capacity to process complex, evolving market information (Wang et al., 2023b).
Quantamental investment refers to combining computer-driven and human-driven research to analyze the amount of market data to construct a portfolio (Tadoori and Guguloth, 2019). For example, alpha factor mining has garnered attention for its ability to identify and exploit market inefficiencies and for its seamless integration with AI-based forecasting methods.
3.6.2. Climate Data for Commodity Trading
Climate data is the records of climate conditions observed at specific locations and times, collected using particular instruments and standardized procedures. Common types of climate data include precipitation, temperature, wind speed, humidity, and satellite imagery of cloud coverage. Climate changes can affect the supply of goods, potentially causing significant price fluctuations and market uncertainty (Stern, 2008; Monasterolo, 2020). By analyzing weather data, investors can better understand and anticipate its impact on financial markets.
3.7. Financial Curriculum and Certificates
Completion of the financial curriculum or certifications requires passing a series of financial examinations. These exam questions contain multimodal financial data, such as textual descriptions, numerical information, graphs, charts, and data tables. Answering these questions needs professional financial knowledge and reasoning capacity. Evaluating the answers of MFFMs on these questions can assess whether these models truly understand financial knowledge (Patel et al., 2024).
4. Multimodal Financial Applications:
Agentic AI Ecosystem
In this section, we first introduce nine financial agents for various typical financial scenarios. Subsequently, we discuss key enablers essential for building an agentic AI ecosystem.
Figure 4. The FinAgent framework powered by FinGPT and Model Context Protocol.
4.1. FinAgents Powered by FinGPT
We envision that AI agents will enable learning systems to take action by observing the complex environment through iterative self-improvement (Durante et al., 2024). This capability could assist in addressing more complex real-world financial tasks. OpenAI [^10] https://cdn.openai.com/business-guides-and-resources/a-practical-guide-to-building-agents.pdf and Google (Wiesinger et al., 2024) recently released detailed guides for agent development, which provide a good starting point for developing FinAgents.
Our SecureFinAI Lab at Columbia has developed several prototypes of FinAgents, powered by FinGPT (Liu et al., 2023, 2024a; Yang et al., 2023; Zhang et al., 2023a; Tian et al., 2024): search agent (Tian et al., 2024), tutor agent (Patel et al., 2024), XBRL agent (Han et al., 2024), and FinRL trading agents (Liu et al., 2018, 2020, 2021, 2022). The search agent can retrieve real-time financial data from the Internet and generate personalized output. The tutor agent teaches professional financial knowledge and complex regulations. The XBRL agent (Han et al., 2024) can analyze SEC filings (in XBRL format) by calling an external retrieval tool and a calculator tool. The FinRL trading agent (Liu et al., 2018, 2020, 2021, 2022) integrates commonly used Deep Reinforcement Learning (DRL) algorithms such as DQN, DDPG, PPO, SAC, A2C, and TD, and facilitates company clients to develop their internal trading strategies.
These agent prototypes follow a standardized development cycle. It enables the development of multiple FinAgents tailored to various financial scenarios. We categorize the nine financial agents into two groups: tool agents and financial service agents.
4.1.1. Tool Agents
  • Search agent: Facing massive multimodal financial data, the MFFM-enhanced FinGPT search agent can retrieve and generate personalized results tailored to the diverse backgrounds and requirements of compound users. These agents would facilitate data-driven decision-making by providing precise, context-aware insights. More importantly, compared to professional financial database platforms such as the Bloomberg Terminal [^11] https://www.bloomberg.com/professional/products/bloomberg-terminal/, the cost of using commercial multimodal large language models (MM-LLMs) (e.g., GPT-4o) or deploying open-source MM-LLMs is lower. Users can easily construct their own customized financial AI search agents, achieving search results that rival those of professional agencies. The effectiveness of such an approach has already been demonstrated by FinGPT search agent (Tian et al., 2024; Zhang et al., 2023b).
  • Tutor agent: There are two recent Guinness World Records (Guinness World Records, 2024): 46,0454604546,04546 , 045 (within $2424$ hours) attendance on an introduction to AI course, and 112,718112718112,718112 , 718 (within $2424$ hours) attendance for a mathematics course. These numbers show a huge demand for online education. MFFMs can provide scalable solutions to meet this demand. For online education platforms, AI tutors equipped with the reasoning capabilities of MFFMs can provide high-quality tutoring services. For students, MFFMs can deliver a personalized learning experience. QFinben (Patel et al., 2024) demonstrates that a pre-trained MFFM model with strong capabilities in undergraduate, graduate, and certificate exams would provide a scalable, personalized solution for AI tutors in business and finance.
  • Robo-advisor: Robo-advisors offer automated, algorithm-driven financial planning and investment management with minimal human intervention. They deliver personalized investment advice and portfolio management to individuals at a lower cost than traditional financial advisors. MFFMs can further enhance Robo-advisors by improving personalized interactions, integrating multimodal data for a comprehensive view of market and portfolio impacts, and providing ongoing adjustments and reminders through continuous user engagement.
  • Coding agent: Coding agents empower investors to rapidly build the personal financial analytical tools (Wermelinger, 2023).
4.1.2. Financial Services Agents
The financial services industry relies on digitized financial information for critical business decisions, such as business operations, investment, and mergers and acquisitions. Digitized financial information includes text, audio, images, and diverse market information. MFFM-powered workflow can integrate diverse multimodal financial data and offer customized financial services tailored to specific needs.
  • Credit scoring agent: Leveraging LLMs, investors could build a credit scoring agent to generate transparent, data-driven credit scores.
  • Auditing agent: In auditing, auditors need to review lots of documents, manage multiple subtasks, and ultimately complete the auditing process. AI agents can autonomously perform complex audit procedures involving tasks such as AI-driven risk assessment or financial statement reviews (Schreyer et al., 2024). By assisting auditors in completing these tasks, AI agents can improve auditing efficiency and reduce human error.
  • Compliance agent: Integrating MFFMs into AI compliance offers organizations a scalable approach to managing regulatory and ethical requirements. Such an integration streamlines compliance workflows, automates complex regulatory analyses, and reinforces ethical AI practices—essential steps for building trust, mitigating risks, and fostering responsible AI developments (Wang et al., 2024a).
  • Report generation agent: Report generation refers to the use of MFFM-powered AI agents to consolidate complex financial data into concise, readable textual content. Regular, accurate, and insightful reports help stakeholders understand performance trends, identify risks, and make informed decisions. MFFM-powered report generation agents enable users to quickly generate high-quality, data-driven, and personalized financial reports.
  • Trading agent: Trading is a complex financial decision-making task influenced by a wide array of market data. MFFM-powered agents can integrate different multimodal market information and output appropriate trading strategies. More importantly, it enables market stakeholders to employ agent systems to get personalized investment suggestions at a low cost. The Financial reinforcement learning (FinRL) trading agent (Liu et al., 2018, 2020, 2021, 2022) offers a user-friendly virtual market environment that includes a wide range of multimodal market data. FinRL integrates commonly used Deep Reinforcement Learning (DRL) algorithms such as DQN, DDPG, PPO, SAC, A2C, and TD, enabling users to develop their own DRL trading strategies. Recently, FinRL 2025 contest [^12] https://finrl-contest.readthedocs.io/en/latest/ proposed the FinRL-DeepSeek project, combining reinforcement learning with LLMs to develop an automated stock trading agent trained on stock price and financial news data. This hybrid approach enhances the capacity to process complex, evolving market information (Wang et al., 2023b).
By summarizing the application scenarios outlined above, it becomes clear that the MFFM-powered agent has great potential to offer market stakeholders scalable, personalized, and cost-effective solutions to multiple complex real-world financial tasks.
4.2. Enablers for Agentic AI Ecosystem
4.2.1. Open Models.
The degree of the open-source model will affect people’s choice and utilization in constructing the agentic AI framework. The openness of a model can be assessed from three dimensions: code, data, and documentation. Many models only open-source a portion of them, and the behavior of misusing the “open source” label is called openwashing (Heimstädt, 2017; Widder et al., 2024; Liesenfeld and Dingemanse, 2024). Openwashing poses a challenge as it introduces confusion into the agentic AI ecosystem.
Model Openness Framework [^13] https://isitopen.ai/ and OpenMDW License  [^14] https://openmdw.ai/ provide the framework to develop an open-source agreement, build an open-source standard, and evaluate the openness of models using a leaderboard (White et al., 2024). This approach would enhance transparency in model usage while preserving the integrity of the entire agentic AI ecosystem.
4.2.2. Agent Protocol
The agent protocol defines standardized communication interaction rules among autonomous agents. It specifies message structures, negotiation mechanisms, and coordination procedures to facilitate efficient collaboration among agents. Currently, there are two commonly used agent protocol:
  • Model Context Protocol (MCP) [^15] https://www.anthropic.com/news/model-context-protocol: MCP is an open standard that enables developers to build secure, two-way connections between their data sources and AI-powered tools.
  • Agent to Agent (A2A) [^16] https://developers.googleblog.com/en/a2a-a-new-era-of-agent-interoperability/: The A2A protocol will allow AI agents to communicate with each other, securely exchange information, and coordinate actions on top of various enterprise platforms or applications.
4.3. Case Study: FinGPT-Powered Search Agent
FinGPT search agent (Tian et al., 2024) can quickly retrieve multimodal financial data customized to the specific needs of individual users or institutional investors and generate personalized content. Fig. 4 provides a generic framework of the FinGPT-powered agents. Interaction begins through a user interface where users input inquiries. Then, the agent will call different MCP clients accordingly to communicate via standard input/output with the corresponding MCP servers.
These MCP servers handle different functions such as: 1) accessing local financial files; 2) searching financial news from remote services like Yahoo Finance through Web APIs; 3) querying market data from databases; 4) analyzing the market sentiment from social platforms like Twitter and Reddit/WallStreetBets. The framework emphasizes user-controlled permissions, explicitly asking for authorization before accessing external data sources, thus maintaining transparency and user privacy. MCP includes local and remote interactions, with remote servers interacting through the networking protocol (HTTP with SSE), ensuring real-time data updates.
5. MFFMs: Progress and Prospect
Recent progress of MultiModal LLMs (MM-LLMs) has attracted research efforts to explore their financial counterparts. First, we present a case study of our first-hand MFFM training experience to illustrate the life-cycle of MFFM development. Then, we summarize the latest development of MM-LLMs. Subsequently, we describe the progress of MFFMs from three aspects: benchmark, model, and dataset. At the end, we highlight the prospects for MFFMs.
Text Tasks Multimodal Tasks Features
Benchmark #Tasks IE TA QA RM FO DM CQA VQA CU NU IU I2T RAG Agent Language
FinBen 24 6 8 3 4 1 1 - - - - - - - EN
Open FinLLM Leaderboard 24 6 8 3 4 1 1 - - - - - - - - EN
QFinBen 1 - - - - - - 1 - - - - - - EN
OmniEval 5 5 - - - - - - - - - - - - EN
InverstorBench 3 - - - - - 3 - - - - - - - EN
FFAMA 1 - - - - - - - 1 - - - - - EN/ZH/FN
MME-Finance 10 - - - - - - - 1 - 3 3 3 - - EN
FinSet-Benchmark 9 1 2 1 1 1 - - 1 1 1 - - - - EN
Table 6. Comparison of financial benchmarks. Text tasks: Information Extraction (IE), Text Analysis (TA), Question Answer (QA), Risk Management (RM), Forecasting (FO), Decision-Making (DM), Complex Question Answer (CQA). Multimodal tasks: Visual Question Answer (VQA), Chart Understanding (CU), Numeral Understanding (NU), Image Understanding (IU), Image-to-Text (I2T).
5.1. Case Study
Figure 5. The lifecycle of model development. It consists of three key stages: (a) Pretraining, where a combination of general and financial corpora is used to pre-train the MFFM, ensuring that it comprehensively understands world financial knowledge; (b) Fine-Tuning, where the pre-trained MFFM fine-tuning on instruction or specific task dataset to enhance its understanding of user intentions or specific tasks; (c) Alignment Tuning, allowing the MFFM to generate content that is more human preference and secure. To make it easier to understand, we have provided examples of data from each stage in (d)–(e).
Fig. 5 illustrates the development life-cycle of an MFFM. The development may consist of three stages: pretraining, fine-tuning, and alignment. Based on our recent open-source projects, we elaborate on each stage of Fig. 5.
(Continual) Pretraining stage: This stage aims to train a model entirely from scratch. A high-quality financial corpus and a powerful base model are the two key ingredients to the success of pre-training an MFFM. Open-FinLLMs (Xie et al., 2024b) employ an 18 billion-token corpus from the general domain and a 52 billion-token corpus from the financial domain. This curated dataset allows the model to keep the general knowledge while getting the financial knowledge. Then, Open-FinLLMs chose Llama3-8B as the base model for the continual pre-training and obtained a financial model called FinLLaMA. The continual pre-training process runs on $6464$ A100 80GB GPUs, approximately $250250$ GPU hours per epoch. FinLLaMA set the maximum sequence length to 8,19281928,1928 , 192 tokens. FinLLaMA surpasses its base model LLaMA3-8B on several financial tasks, highlighting the effectiveness of (continual) pretraining.
Fine-tuning stage: This step aims to enable the model’s multimodal capabilities, enhance the model’s instruction-following capabilities, and optimize performance on downstream financial tasks. Building upon FinLLaMA, its multimodal extension, FinLLaVA addresses multimodal financial tasks by employing multimodal instruction tuning. The instruction-tuning dataset comprises $1.431.43$ million image-text pairs. Instruction tuning is conducted on eight NVIDIA HGX H20 80GB GPUs, with the entire process requiring approximately 30 hours for one epoch. FinLLaVA outperforms all open-source MM-LLMs chart understanding tasks and is second only to the closed-source GPT-family MM-LLMs.
Alignment stage: This step aims to guide fine-tuned MFFMs to generate human-preferred and safety output. FinTral (Bhatia et al., 2024) includes an alignment tuning process. First, an alignment dataset is constructed. FinTral fed the instruction dataset into both a high-capacity LLM, such as GPT-4, and a less capable casual LLM. The output from the high-capacity LLM is labeled as positive samples, while those from the casual LLM are labeled as negative samples. Then, alignment tuning is conducted using the Direct Preference Optimization (DPO) method (Rafailov et al., 2023). After alignment tuning, FinTral not only generates the output that is aligned with human preferences but also greatly reduces the hallucinatory content.
5.2. Progress of MFFMs
We reviewed the progress of MFFMs from three aspects: existing MFFMs’ benchmarks, the development of MFFMs, and multimodal financial datasets.
5.2.1. Benchmarking MFFMs’ Performance
Measuring Performance on various financial tasks is crucial to understanding their capabilities quantitatively. Currently, multiple financial benchmarks provide comparisons from different perspectives:
  • FinBen (Xie et al., 2024a): IT includes $4646$ datasets spanning $2424$ financial tasks and covers seven critical tasks: information extraction (IE), textual analysis, question answering (QA), text generation (TG), risk management (RM), forecasting (FO), and decision-making (DM). FinBen evaluated 30 representative LLMs and identified several key findings: LLM performed well in IE, and text analysis, but its performance in complex tasks such as high-level reasoning and text generation and prediction still needs to be improved.
  • Open FinLLM leaderboard (Lin et al., 2024a): Building on FinBen (Xie et al., 2024a) and PIXIU (Xie et al., 2023), this leaderboard aims to maintain an open platform that encourages innovative adoption and improved model. Open FinLLM Leaderboard provides an interface between academia, the open-source community, the financial industry, and other stakeholders. Open FinLLM Leaderboard creates a collaborative and open ecosystem by continuously updating new datasets, tasks, and model performance.
  • High-Quality financial benchmark (QFinBen) (Patel et al., 2024): QFinBen explores the reasoning capabilities of LLM in complex financial questions. QFinBen assembled a dataset of 8,050 questions sourced from undergraduate and graduate finance, accounting, and economics examinations alongside professional financial exams such as the CFA, CPA, and FRM. QFinBen tests the dataset by using the GPT-4o, Llamma 3.1-405B, and Mistral Large 2. The findings indicate that LLMs still struggle to pass these complex examinations, highlighting their current limitations in addressing sophisticated financial reasoning challenges.
  • FinSet-Benchmark (Bhatia et al., 2024): It’s part of FinTral (Bhatia et al., 2024), containing 13 LLMs on seven text-based financial tasks, and 9 $M$-LLMs on Chart Understanding.
  • MME-Finance (Gan et al., 2024): MME-Finance is a bilingual financial visual question and answer (VQA) benchmark. MME-Finance conducted extensive experimental evaluations on 19 MM-LLMs to test their perception, reasoning, and cognitive abilities on financial multimodal data. The results show that MM-LLMs that perform well in general benchmark tests may perform poorly on MME-Finance. Specifically, these models show poor performance in understanding candle charts and technical indicator charts.
  • FFAMA (Xue et al., 2024): It’s an open-source benchmark for financial multilingual multimodal question answering (QA). It includes 1,758 meticulously collected question-answer pairs from university textbooks and exams, spanning 8 major subfields in finance, including corporate finance, asset management, and financial engineering. FFAMA assesses a variety of SOTA MM-LLMs, revealing that FAMMA presents a considerable challenge for these MM-LLMs. Even advanced models such as GPT-4o and Claude35-Sonnet attain only a 42% accuracy.
  • OmniEval (Wang et al., 2024c): It is the first RAG benchmark in the financial domain. OmniEval evaluates the RAG framework from a multi-dimensional that includes (1) a matrix-based RAG evaluation system that classifies queries into five tasks and 16 financial topics, thereby structuring the evaluation of different query scenarios; (2) A multi-stage evaluation system that evaluates search and generation performance to evaluate the RAG process comprehensively. (3) Robust evaluation metrics derived from rule-based and LLM-based evaluation metrics. The results of OmniEval highlight that the RAG can effectively integrate external knowledge to improve the accuracy of the generated results in a variety of tasks. However, the evaluations also reveal that the RAG system struggles with complex multi-hop reasoning and numerical understanding.
  • InverstorBench (Li et al., 2024): It’s the first LLM-based financial agent benchmark. InverstorBench provides a comprehensive performance evaluation of 13 different LLMs across varied market scenarios, including stock trading, cryptocurrency trading, and ETH trading. This benchmark shows that proprietary models (e.g. GPT-4) generally exhibit better financial decision-making capabilities under complex market conditions. However, InverstorBench also indicates that the performance of different LLMs varies in stock, cryptocurrency, and ETF trading. This variability not only underscores the inherent complexity of financial markets but also emphasizes the critical importance of model selection and fine-tuning on specific financial corpus.
These benchmarks provide an overview of the current landscape of applying LLM to financial tasks. We can find that: 1) LLMs/MM-LLMs can effectively improve the capabilities of information extraction relevant tasks and basic financial text analysis. Such improvements can help users build automated financial data processing systems, thus saving manual efforts and reducing human errors. 2) Current LLMs/MM-LLMs still have limitations in their capacity to answer complex financial questions, comprehend numerical values, and interpret charts and tables. This underscores the urgency of developing MFFMs tailored for financial multimodal data.
5.2.2. MFFMs
Typically, MFFMs are built from open-source LLMs, which serve as the backbone. These MFFMs are pre-trained and fine-tuned using the specialized financial dataset. To more clearly illustrate the existing MFFMs, we draw the development path in figure LABEL:fig:modeltree. We aim to provide readers with a comprehensive understanding of the advancements in Multimodal Foundation Financial Models. We highlight representative MFFMs in below:
  • Open-FinLLMs (Xie et al., 2024b): We have introduced Open-FinLLMs in Subsection 5.1. Open-FinLLMs consist of two Financial LLMs: FinLLaMA and FinLLaVA. The experiment results demonstrate that FinLLaMA gets superior performance over LLaMA3-8B, LLaMA3.1-8B, and BloombergGPT in text classification, credit scoring, fraud detection, Q&A, Sentiment Analysis, NER, and decision-making tasks. FinLLaVA outperforms GPT4 and other Financial LLMs in understanding tables and charts. The results from Open-FinLLMs highlight the effectiveness of training financial domain-specific LLMs/MM-LLMs.
  • FinTral (Bhatia et al., 2024): It is a suite of state-of-the-art MFFMs built upon the Mistral-7B model and tailored for pure text and multimodal financial analysis. FinTral is pre-trained on 20 billion tokens of domain-specific data in the first step, followed by instruction fine-tuning and alignment with AI feedback. Subsequently, FinTral gets further instruction fine-tuning on multimodal instruction data. FinTral demonstrates good zero-shot capabilities, outperforming GPT-4 in five of eight text-based tasks. Moreover, FinTral’s multimodal performance surpasses that of all other open-source MM-LLMs, ranking just behind GPT-4V.
  • FinVis-GPT (Wang et al., 2023a): It’s a new MFFM specialized in financial chart analysis. FinVis-GPT is pre-trained on a finance-oriented alignment/instructing following dataset, which includes various types of financial charts and their corresponding descriptions. The experiment results show that FinVis-GPT can interpret financial charts and provide valuable analysis.
These studies have demonstrated that MFFMs already take important roles in multiple financial tasks. These works lay the foundation for more sophisticated applications of AI in finance, potentially transforming the landscape of financial analysis. Although the performance of these MFFMs in some complex financial tasks still needs to be improved, these findings also highlight the significant potential for future development of MFFMs. Future work will focus on further expanding the applicability of MFFMs in more complex tasks and diverse financial scenarios.
5.2.3. Multimodal Financial Datasets
Different stages rely on different types of training data. A high-quality training dataset will affect the capacity of the trained MFFMs. This part will discuss each stage’s dataset construction and characteristics.
  • Pre-training dataset. As the first stage, pre-training data aims to provide multimodal financial knowledge for models and enable the model to align the different modalities. During this stage, different models curate their unique training corpora. A representative training dataset is BloombergGPT’s FinPile (Wu et al., 2023). It comprises a total of 345 billion tokens from public data and 363 billion tokens from proprietary data.
  • Instruction-tuning dataset. This stage aims to teach models to better understand the instructions from the demanded tasks to boost zero-shot capacity. OpenFinLLaVA first assembled a comprehensive multimodal dataset, subsequently utilizing GPT-4o to extract financial content selectively. Ultimately, OpenFinLLaVA created an extensive collection of 662k multimodal pre-trained datasets, comprising images, charts, and tables. FinVis-GPT utilized historical daily data from Chinese A-share stocks to create visualizations, distributing the output into 80% candlestick and 20% line charts. FinTral utilizes several datasets to build a visual pretraining dataset. Additionally, the Llava Instruct dataset is employed to enhance instruction understanding in the multimodal LLMs, resulting in the creation of the instruction tuning dataset, FinVis-IT.
5.3. Prospects of MFFMs
5.3.1. Multimodal retrieval-augment generation (MRAG)
The ability to retrieve relevant information efficiently from a large database is crucial for the success of financial AI systems. Enhancing retrieval-augmented generation capabilities will enable more precise and contextually aware responses from AI models, significantly improving their usefulness in complex financial decision-making processes.
5.3.2. Fine-tuning and quantization methods
For general-purpose LLMs to be effective in finance, they need to be fine-tuned with domain knowledge that captures the nuances of financial markets and instruments. Additionally, model quantization should be considered to optimize inference performance in terms of speed and resource consumption, ensuring that the models can be deployed effectively in real-time environments. FinGPT-HPC (Liu et al., 2024a) and FinLoRA (Wang et al., 2024b) are two examples of applying quantization techniques in the fine-tuning process.
5.3.3. Customizing pretrained models to use senarios
Customizing pre-trained models to cater to specific use cases can significantly enhance their effectiveness. For instance, a model trained for general customer service may require adjustments to handle complex investment queries or to comply with specific regulatory requirements in finance.
6. Challenges and Opportunities: Towards Financial AI Readiness and Governance
Adapting MFFMs to real-world financial scenarios will face several challenges. This also presents research opportunities.
6.1. Proprietary Multimodal Financial Data
Proprietary data is important for financial analysis and decision-making because it provides unique insights: (BloombergGPT - Proprietary data)
  • Internal trading data: The financial institutions have the capability to track and analyze their transaction data, offering insights into behavioral patterns and market trends.
  • Credit scoring data: Financial entities possess data regarding the credit histories of individuals and corporations, which is essential for risk management.
  • Market research data: Data gathered through specialized market research or customer feedback can aid financial firms in understanding consumer demands and market dynamics.
  • Real-time streaming data: Certain institutions have access to real-time transaction flow data, which significantly facilitates algorithmic trading.
  • Private financial reports: Some companies may have access to confidential financial information about partners or potential investment targets.
  • Proprietary economic indicators: Large institutions may develop their own macroeconomic or microeconomic indicators based on exclusive datasets and analyses.
  • Alternative data: This includes satellite imagery, mobile app data, and social media activities, which can provide additional perspectives and information for investment decisions.
Synthetic Multimodal Data. The training of MFFMs has two challenges: 1) Data privacy - the sensitivity of financial data limits its use in constructing training datasets; 2) Data quality - a scarcity of high-quality multimodal financial data, with the existing data mainly consisting of ¡Chart Image - Text¿ pairs that lack balanced representation from various modalities. These challenges constrain the further development of MFFMs’ capabilities. Therefore, enhancing the diversity and quality of multimodal financial data has become a critical need. Synthetic Multimodal Data provides a potential solution to these issues.
Synthetic data (Assefa et al., 2020) is from a generative process that learns the properties of real data but cannot be traced back to the raw data sources. The objective of synthesizing multimodal data is to generate data that accurately reflects the real distribution while also ensuring it cannot be traced back to the original sources to fulfill privacy requirements. There have been multiple demos in the medical field that have used synthetic multimodal data to augment datasets, which demonstrates the effectiveness of synthesizing multimodal data (Wendland et al., 2022; Pezoulas et al., 2024). However, in the financial domain, Potluru et al. (2023) provides a comprehensive review of the field of financial data synthesis and points out the current lack of efficient multimodal data synthesis methods. This highlights the challenges and opportunities of synthetic multimodal financial data.
Figure 6. Guardrail framework for FinGPT-powered agents using Zero-Knowledge Proofs (ZKPs) and blockchain technologies.
6.2. Digital Regulatory Reporting (DRR)
A chatbot with multimodal capabilities (Wang et al., 2024a)(Tian et al., 2024) may help automate the financial regulatory process. For example, when lawyers perform case studies, chatbots can quickly search and summarize relevant legal provisions and historical cases, saving time over manual searches. When accountants prepare financial statements, chatbots can assist in checking compliance with generally accepted accounting principles (GAAP).
However, the financial regulatory landscape presents unique challenges to MFFMs. First, the complex framework and overlapping jurisdictions of financial regulation make the compliance process complex. In the European Union (EU), the European Supervisory Authorities (ESAs) need to collaborate closely with national regulators to maintain a cohesive regulatory environment across Member States  (Directorate-General for Financial Stability, Financial Services and Capital Markets Union, 2022; Cherednychenko, 2021). The U.S. financial regulatory framework is fragmented, comprising federal and state laws. It involves various entities, including federal agencies, state regulators, interagency bodies, and international regulatory fora, with overlapping jurisdictions  (Labonte, 2023). Second, financial regulation requires processing multimodal data from different sources. This includes structured data, such as SQL databases and XBRL filings; unstructured data, such as regulatory texts; dynamic and noisy transaction data; and code in financial product management systems. The format and complexity of each data type vary greatly, which creates a challenging environment for AI compliance.
XBRL: eXtensible Business Reporting Language (XBRL) is an open international standard for business reporting, in order to streamline financial data creation, dissemination, and analysis. XBRL facilitates information exchange among investors, regulatory bodies, and market participants, boosting market transparency and regulatory compliance. Over the last two decades, most global economies have adopted XBRL for financial information sharing. However, the complexity of XBRL necessitates specialized knowledge for proper understanding and analysis, posing a steep learning curve for businesses and a challenge for widespread accessibility by the general public.
An XBRL agent (Han et al., 2024) will simplify data aggregation and support informed decision-making. It may provide users with easy access to financial intelligence. How we interact with financial data is no longer the exclusive domain of a few individual experts but a valuable resource for everyone. On the FinanceBench dataset, a public dataset comprising SEC document-related questions (150 openly available sample questions), (Han et al., 2024) evaluated the current AI ChatBots (e.g., ChatGPT, LLama2, FinGPT). The results show that its accuracy in answering financial questions is only about 19%   30%, which is far from the professional level. The errors may come from several aspects: 1). Ambiguity in complex financial terminology; 2). Errors in interpreting and extracting data from financial documents; 3). Calculation errors (e.g., financial ratio calculation and aggregation).
Common Domain Model (CDM): The Common Domain Model (CDM), a standardized, machine-readable/machine-executable data and process model for multiple financial products, is a promising fundamental solution to address the above challenges. Developing a CDM for XBRL using the Multimodal Large Language Model can handle various document formats, including PDFs, scanned documents, and webpages. High-quality document reading can effectively reduce errors during document reading and support financial documents in diverse scenarios. Furthermore, using MM-LLMs as the backbone and combining multiple external tools or RAG techniques to construct a standard agentic workflow can mitigate ambiguity in financial terminology and numerical calculation errors during format conversion.
6.3. Ethical Challenges
There are intensified ethical concerns with MFFMs. Mishandling sensitive information and thus making unfair, biased judgments can be disastrous to financial institutions. Analysts who trust flawed MFFMs will make bad investment decisions and improper risk assessments. Small missteps can cause significant client dissatisfaction and negative media attention.
Persistent ethical issues include:
  • Security and privacy: It is vital that FinLLMs have airtight security to prevent leakage of sensitive information. Example: Samsung employees accidentally leaked company secrets when prompting ChatGPT for help.
  • Copyright infringement: FinLLMs trained on Internet data are not allowed to output copyrighted data to end users. Example: The New York Times sued OpenAI and Microsoft for using millions of its articles; Perplexity was accused of using articles from The Wall Street Journal or The New York Post to populate its RAG database and generate responses to user queries.
  • Systematic bias: In decision-making processes, FinLLMs’ systematic bias may lead to unfair discrimination towards certain racial groups. According to Zillow and Consumer Reports, LLMs may quote African Americans at higher prices in home mortgages and car insurance due to historical segregation towards disaster-prone areas.
  • Transparency, explainability, and accountability: It is important to ensure that FinLLMs are transparent, explainable, and accountable, providing clear responses, especially in finance where every decision has significant implications. J.P. Morgan Chase established its firmwide Explainable AI Center of Excellence (XAI COE) for research on explainability and fairness in finance.
Newly-emerging ethical issues include:
  • Truthfulness: LLMs consistently hallucinate, creating false statements. In business and finance, hallucinations are problematic because LLMs’ output must exactly match information extracted from earnings reports when queried. Microsoft faced backlash when Bing AI hallucinated when analyzing Gap and Lululemon’s earnings reports during a demo.
  • Sycophancy: LLMs demonstrate sycophancy, catering their outputs to match user beliefs rather than being truthful. Sycophancy is problematic when it causes inaccurate confirmation of financial analysts’ and accountants’ math.
  • Compliance with professional norms: LLM responses must follow professional norms to avoid implicit toxicity in training data. This is vital to preserve company culture and public relations.
  • Law and regulatory compliance: FinLLMs must comply with current financial laws and regulations when making decisions and chatting with end users. According to the Consumer Financial Protection Bureau, FinLLMs must comply with regulations in operations like fraud detection, citing concerns like discrimination against minority racial groups.
6.4. Misinformation and Hallucination
In the financial domain, the accuracy of information is important for the integrity of market operations, risk management, compliance, and financial decisions. There are two sources of inaccurate financial information: dissemination of misinformation and hallucination from the model’s output.
Misinformation is from various media channels (Rangapur et al., 2023) and the misuse of LLMs to generate misinformation (Zhou et al., 2023; Cao et al., 2024b). Detecting financial misinformation is a challenge. To address this issue, FMDLlama (Liu et al., 2024b) fine-tuned the LLaMA-3 model on the Fin-Fact dataset (Rangapur et al., 2023) to detect financial misinformation. This case presents a feasible solution. By leveraging LLMs, an agentic framework can be developed to detect dynamically evolving financial misinformation.
Hallucination is factually incorrect output from LLMs due to their training on vast and diverse datasets. Ensuring the accuracy and reliability of LLM-generated outputs is crucial for their application in the financial industry. Kang and Liu (2023) quantified financial hallucinations and explored several potential solutions to mitigate them, including few-shot learning, decoding by contrasting layers, and RAG.
6.5. Guardrail Framework for FinAgents
Wide adoption of financial agents raises concerns about privacy, security, and trust. We first outline the potential risks inherent in financial agent workflows. Then, we propose a guardrail framework that leverages zero-knowledge proofs (ZKPs) and blockchain technology. Blockchain and ZKPs ensure that financial agents’ actions remain secure, verifiable, and immutable, fostering transparency and trust.
6.5.1. Risk in financial agents workflow.
Three FinGPT-powered agent prototypes were introduced in Section 4.1 following an agentic pipeline: An agent calls various tools via the Model MCP to retrieve relevant content from local files, backend databases, remote models, and the Internet. We identify several major risks across different financial scenarios (red points in Fig. 6), referring to Linux’s AI Readiness Governance Framework:
  • Information leaked to host model. Enterprise users may frequently employ FinGPT search agent to process local files (e.g., income sheet) for compliance tasks like internal audits, risk assessment, or regulatory reporting. These files contain personally identifiable information or commercially sensitive data. Corporate employees and students increasingly rely on AI tutors for financial knowledge from local textbooks. These books usually have copyright restrictions. During multi-round dialogues, the sensitive local files or copyrighted content may leak to external models (red line and point in Fig. 6).
  • Misinformation. Users employ FinGPT search agent to obtain real-time information from financial websites and social media. However, the generated responses may contain misinformation and biased content (orange line and point in Fig. 6).
  • Lack of source data access control. FinGPT-powered agent could access external data sources (e.g., subscription-based websites). However, since these sources may enforce different access control policies, users might inadvertently access data that they are not authorized to retrieve directly from the original source. This unauthorized access may also lead to copyright issues (orange line and point in Fig. 6).
  • Hallucination. LLM-based output may contain hallucination content, which refers to information that appears plausible but is factually incorrect. Inaccurate output can lead to costly errors, operational inefficiencies, and misinformed decisions. (point in Fig. 6).
To mitigate these identified risks, integrating Zero-Knowledge Proof (ZKP) protocols and blockchain technologies represents a promising solution. These technologies form the foundation of a novel guardrail framework designed specifically for financial agents, as illustrated in Fig. 6 (blue line, points - ).
6.5.2. Zero Knowledge Proofs (ZKPs) for Privacy-Preserving
Zero-Knowledge Proofs (ZKPs) are cryptographic protocols that let a prover convince a verifier of a statement’s correctness without disclosing any underlying secrets. The ZKPs ensure three key properties: completeness (an honest execution always produces a verifiable proof), soundness (no one can forge a valid proof without performing the correct computation), and zero-knowledge (the verifier learns nothing beyond the truth of the statement). zkLLM (Sun et al., 2024) has demonstrated that ZKP protocols help protect the privacy of the large language model parameters (usually considered as intellectual property of model producers). For LLMs with 13B parameters, zkLLM can verify the inference process in less than 15 minutes, and the generated proof file has less than 200 KB.
To ensure privacy in the agent workflow, the agent generates a ZKP proof file and uploads it to the blockchain (blue line ), demonstrating that the actions (search steps, inference steps, and output procedure) strictly adhere to pre-established inference schemes without exposing sensitive or proprietary data to remote model (red line and point ). To enhance access control for external source data, the external participants generate a ZKP file for copyright (blue line ). When the local agent takes actions, copyright permissions are granted by the blockchain, preventing unauthorized retrieval of external content (orange line and point ).
6.5.3. Blockchain-Layered Agent Life Cycle.
The generated ZKP protocol files, agent updates, copyright policies, and regulatory documents are recorded on a permissioned blockchain (e.g., Hyperledger Fabric or Corda). Participants interact with the blockchain by submitting cryptographic hashes referencing agent updates, inference steps, or compliance logs (blue line - ). Agents update trusted source lists (blue line ) to avoid misinformation from external content (orange line, point ) and load inference schemes (blue line ) to prevent hallucinations (point ). This creates an immutable audit log, enabling stakeholders to verify agent compliance with approved procedures.
These two components jointly enable FinGPT-powered agents to incorporate local data securely and produce on-chain verifiable proofs of correctness and compliance. By preserving confidentiality of sensitive data (via ZKPs) while anchoring essential references in a tamper-proof ledger (via blockchain), our approach harmonizes the conflicting needs of safety, confidentiality, transparency, and regulatory oversight.
7. Discussion and Conclusion
This paper offers a comprehensive overview of Multimodal Financial Foundation Models (MFFMs), highlighting their state of readiness. First, we review the multimodal financial data and application scenarios. Then, we describe the progress and future prospects of MFFMs. We further analyze the challenges and opportunities faced by MFFMs to achieve AI readiness.
By summarizing the current state of readiness, multimodal financial application scenarios, multimodal financial data, and the development of MFFMs, this paper aims to inspire future research and innovation in both the academic and financial industries.
As we navigate the integration of machine learning in business and finance, it is paramount to address the multifaceted challenges that arise from the unique characteristics of multimodal financial data and the new capabilities of MFFMs. Here, we outline strategic directions and considerations that will enhance the financial AI readiness for individuals and institutions:
  • Multilingual and multimodal. Financial data is inherently complex, often presented in various modes, including text, numerical data, images, and more. An effective financial AI framework must be capable of interpreting and integrating these diverse multimodal data seamlessly. Furthermore, the global nature of finance demands multilingual capabilities to ensure that insights can be gleaned from data across different languages and regions. AI models should be equipped to handle multiple tasks simultaneously, such as risk assessment, fraud detection, and customer service, to provide comprehensive solutions.
  • Open datasets and question sets. Open datasets will facilitate the training of the more Powerful MFFMs. Adding complex open financial questions into training datasets can further enhance their reasoning capabilities. Furthermore, Public open datasets and question sets assist in establishing a standard benchmark for evaluating MFFMs.
  • Open leaderboard of MFFMs and FinAgents. Building an open leaderboard enables rapid evaluation of the progress and characteristics of different MFFMs. It will facilitate the development of an agentic AI ecosystem.
  • Blockchain. Data privacy and protection of model intellectual property are the challenges when developing an agentic AI ecosystem. Blockchain technology allows multiple organizations to collaboratively train a shared model while safeguarding data privacy, preventing leakage of model parameters, and transparently verifying each participant’s contributions.
Acknowledgements.
The authors thank Keyi Wang for helping draw Fig. 4. Xiao-Yang Liu Yanglet acknowledges the support from Columbia’s SIRS and STAR Program, The Tang Family Fund for Research Innovations in FinTech, Engineering, and Business Operations. Xiao-Yang Liu Yanglet also acknowledges the support from the NSF IUCRC CRAFT Center research grant (CRAFT Grant 22017) for this research. The opinions expressed in this publication do not necessarily represent the views of NSF IUCRC CRAFT.